### v3

@@ 1 Introduction
However, the way the NP surface influences the protein structural changes within the corona remains debated [15, 16, 17, 18]. For instance, theory predicts that proteins near a hydrophobic interface should unfold and aggregate less than those in bulk, and that increasing temperature does not facilitate unfolding and reduces aggregation [19]. Conversely, on a hydrophilic NP, the prediction is that proteins adsorb while remaining folded, and as temperature increases, they unfold but do not desorb [20].
Here, by combining different experimental techniques with Machine Learning (ML) analysis, we characterize the structural changes in Fib over temperature when exposed to carbon NPs (CNP) and silicon dioxide (SiO2) NPs. We use monodisperse nanoparticles of similar size but differing surface chemistry. CNPs and SiO2NP have a comparable geometrical and hydrodynamic diameter to prevent any effects caused by a differing surface curvature. Both exhibit a negative zeta potential at neutral pH [5, 21], indicating the presence of weakly acidic groups on the surface. In the case of CNPs, these species include acidic carboxylic or phenolic groups formed during synthesis, while for SiO2NP, they consist of surface hydroxyl groups (silanols, Si-OH). The density of acidic groups is expected to be in the same range for both samples (around 4-6 groups per nm2) [5]. However, the two surfaces exhibit significant differences. While in CNPs the oxygenated moieties are bound to a hydrophobic surface composed of C-C bonds, in the case of silica, silanols are attached to a more hydrophilic framework made up of polar Si-O-Si bridges. Therefore, while SiO2NPs are hydrophilic, CNPs are more hydrophobic.
We employ several spectroscopic techniques: UltraViolet Resonance Raman (UVRR) spectroscopy [22], Circular Dichroism (CD) [23], and UV absorption [24] spectroscopy.
All these techniques are non-destructive and do not require chemical labeling. For instance, by setting the excitation wavelength to correspond with the electronic transition energy of the aromatic amino acids around 220-230 nm, the Raman signals associated with tryptophan (Trp) and tyrosine (Tyr) are selectively enhanced, providing insights into the protein’s tertiary structure and how its chromophores interact with their surroundings [25, 26]. On the other hand, CD spectroscopy, which is based on the differential absorption of left- and right-hand circularly polarized light, can indicate the presence of disorder in the protein’s secondary structure [27, 28]. Lastly, UV absorption spectra are very sensitive to the presence of aggregates and precipitates, which can elastically scatter light according to Rayleigh’s law [22].
We analyze our high-dimensional, multi-component data from various spectroscopic techniques using a novel unsupervised ML workflow designed to quantify and visualize how protein conformational changes occur as temperature varies. The novelty of our proposed ML model rests on two key algorithms: the noisy-resistant Wasserstein distance [29, 30] and t-distributed stochastic neighbor embedding (t-SNE) [31, 32, 33].
These statistical tools are particularly advantageous because both remain unaffected by the curse of dimensionality, i.e., the challenges and limitations of working with high-dimensional data, such as overfitting, sparse distribution, computational complexity, and saturation of distances [34, 35, 36]. Indeed, high-dimensional data render standard similarity tools, such as the Euclidean and Manhattan distances commonly used in cheminformatics [37], ineffective (Section 1 in Electronic Supplementary Information, ESI).
Furthermore, t-SNE successfully replaces principal component analysis (PCA) [38, 39, 34], which is routinely employed to visualize spectral data in two-dimensional space. PCA cannot effectively handle data from diverse sources, mainly because, by preserving large distances arising from different spectra, it misses interesting structures, such as clustering, at smaller distances [^1] This is why PCA is susceptible to outliers in the data (Section 2 of ESI). In contrast, t-SNE overcomes this issue. When PCA is used as an initialization step, t-SNE can preserve both global and local data structures, yielding a meaningful embedding [33, 40].
Our multi-component data are organized into vectors with 856 dimensions, measured between 22 and 86C at temperature intervals of $\Delta T=4^{\circ}$. Overall, our analysis reveals that the protein structural disorder gradually increases for Fib adsorbed onto CNPs as the temperature $T$ rises, suggesting a ’frozen’ partially unfolded structure that is only weakly dependent on $T$. In contrast, the structural behavior for Fib adsorbed onto SiO2NPs is similar to that of free Fib, both characterized by a significant increase in disorder around 50C. The t-SNE cluster analysis further supports these findings. For free Fib, the substantial increase in disorder is experimentally understood as a direct result of Fib aggregation. Therefore, our novel ML analysis of multi-component spectral data offers an experimental confirmation of theoretical predictions, providing insights into the interaction mechanisms between proteins and NPs. It highlights the importance of NP surface properties in protein adsorption and stability, and represents a new and solid method to analyze the controversial structural changes during biomolecular corona evolution.
Figure 2: Physicochemical characterization of the NPs and the Fib corona. (a) Transmission Electron Microscopy (TEM) image of CNPs with a nominal diameter of $\sim$120 nm. After image processing, they were calculated to have a core diameter of (118.3 $\pm$ 40.8) nm. (b) DLS distribution for pristine (‘plain’) CNPs and in solution with Fib at WR between 0.1 and 1.0. In all cases, the polydispersity index (PDI) is less than 0.12. (c) DCS analysis for the corresponding cases. (d) TEM image of SiO2NPs with a nominal diameter of 100 nm. Their calculated size of (95.4 $\pm$ 7.8) nm. (e) DLS of SiO2NPs with PDI of 0.01, and (f) The corresponding DCS analysis.
2 Results
2.1 Experimental results
2.1.1 NP characterisation by Dynamic Light Scattering and Transmission Electron Microscopy and Fib corona stability
We follow the experimental protocols developed in our lab to obtain the protein corona formed on top of carbon and silica NPs [41]. We first characterize the pristine NPs by Transmission Electron Microscope (TEM). By using an electron beam on dry samples, TEM provides high-resolution images and characterizes the size populations of NPs (Fig.2a and d). To study the samples in the colloidal suspension, we use Dynamic Light Scattering (DLS). In DLS, a laser tracks the movement of NPs due to Brownian diffusion, allowing the measurement of the autocorrelation function and diffusion coefficient, which in turn determine the particles’ hydrodynamic size distribution and polydispersity index (PDI) directly on the colloidal suspension. We find that both CNPs and SiO2NPs are monodisperse in aqueous solution and stable against aggregation (Fig.2a and d).
Next, we check whether the solutions of NPs interacting with Fib are stable at room temperature. In fact, it is well known that the addition of proteins in solution can induce NP aggregation [42] or disaggregation [43], and aggregation in simple media can occur at much lower protein concentrations than in complex media with many components [44].
Therefore, we use various techniques to analyze the stability of NP-protein corona complexes against aggregation after incubation at different protein-to-NP weight ratios (WRs), ranging from 0.1 to 1.0. By employing DLS and Differential Centrifugal Sedimentation (DCS), we do not observe aggregates for CNPs at the tested WRs (Fig.2 b and c). However, for the same range of WRs, aggregation systematically occurs with SiO2NP (Fig.2 e and f).
Therefore, the difference in the NPs’ chemistry likely impacts protein adsorption and aggregative propensity differently. Next, we set our experimental conditions to WR=1 and compare the cases of both CNP (colloidally stable) and SiO2NP (colloidally unstable) with a free Fib solution at various temperatures.
First, we observe, by TEM characterization, that Fib at blood-like concentration agglomerates when the temperature reaches 88C. The aggregates clearly resemble the structure of fibrin fibers [45] (Fig.3).
Next, our DLS analysis indicates that the hydrodynamic diameter of free Fib and Fib with silica NPs varies with temperature. In both cases, there is an elevated PDI¿0.4 corresponding to multimodal distributions (Fig.2 e and f), with aggregation occurring at 52C and above.
In contrast, the sample containing Fib and carbon NPs shows a low PDI¡0.2, associated with an unimodal distribution, at any temperature (Fig.2 b and c). However, at 52C, it exhibits a shift in PDI that may be related to a gradual denaturation of Fib, which affects the solution’s colloidal stability.
2.1.2 UV Resonance Raman Spectroscopy
UVRR spectroscopy is a powerful tool for detecting structural changes in organic biomolecules, such as proteins and peptides, without the need for any chemical label [46]. Compared to spontaneous Raman spectroscopy using visible excitation sources, the UVRR technique offers undeniable advantages. The visible Raman profiles of proteins typically result in a complex superposition of several signals, making analysis difficult. Thanks to the selective resonance enhancement achieved by appropriately tuning the excitation wavelength in the deep UV range (200-300 nm), UVRR spectra of proteins can be simplified, and the signals arising from selected chromophores or molecular portions of the molecule can be more easily disentangled [47, 48, 49]. In this regard, using Synchrotron Radiation (SR) as an excitation source for UVRR experiments provides the added advantage of fine-tuning the excitation wavelength to select the best working conditions for the system under investigation [50, 51, 52]. Additionally, the high sensitivity of the UVRR technique is crucial for investigating proteins in more diluted conditions compared to visible Raman spectroscopy while maintaining satisfactory quality in the spectra. This, in turn, provides the opportunity to simultaneously extract information from the UVRR spectra of proteins in aqueous solutions, both from the Raman signals attributed to the polypeptide solute and the water molecules in the hydration shell [53, 51].
We collect the UVRR spectra for Fib in PBS buffer, both in the presence and absence of NPs, using 226 nm as the excitation source (Fig. 4 a-c). This excitation energy falls within the absorption range of the aromatic rings of Tyrosine (Tyr) and Tryptophan (Trp) residues, resulting in UVRR spectra of Fib that are predominantly characterized by the vibrational signals of these aromatic amino acids [47, 54] (Fig. 4 a).
In the Raman spectrum of the Fib protein, we can identify several key resonance Raman features associated with the Trp vibrational modes, including signals at 1012 $\mathrm{cm^{-1}}$ (Trp W16), which are attributed to the symmetric benzene/pyrrole out-of-phase breathing, at 1340-1360 $\mathrm{cm^{-1}}$ (Trp W7), which is attributed to the Fermi resonance between N1-C8 stretching in the pyrrole ring and combination bands of out-of-plane bending, and at 1554 $\mathrm{cm^{-1}}$ (Trp W3), which relates to the C2-C3 stretching mode of the pyrrole ring [55]. In the same spectral range of 1000-1800 $\mathrm{cm^{-1}}$, Tyr Raman bands are also detected, specifically the signals at 1176 $\mathrm{cm^{-1}}$ (Tyr Y9a), which are associated with in-plane C-H bending, and at 1616 $\mathrm{cm^{-1}}$ (Tyr 8a), which is due to the in-plane ring stretching of Tyr [56]. Additionally, a signal at approximately 1770 $\mathrm{cm^{-1}}$ is observed in the UVRR spectrum of Fib, likely corresponding to a C=O stretching vibration.
Figure 4: UVRR spectra for free Fib, and Fib with NPs. The 226 nm-excited UVRR spectra are collected for Fib 2mg/mL in PBS pH 7.4 (a), Fib with CNPs (b), and Fib with SiO2NPs (c) as a function of temperature, reported in the two spectral ranges 1000-1800 and 2800-3800 $\mathrm{cm^{-1}}$. The main vibrational features in the spectrum of Fib are labelled in the panel (a). Temperature range and steps are indicated in the colour scale at the right. Dashed lines in panels (b) and (c) represent the spectra of Fib alone at the lowest temperature of 22C.
Since the aromatic rings of Trp and Tyr can participate in hydrophobic, cation-$\mathrm{\pi}$, and hydrogen-bond interactions with nearby residues or surrounding solvent molecules [55], the spectral feature parameters of the Raman bands associated with these residues provide valuable insights into their local environment. For instance, it has been reported that the vibrations of aromatic ring side chains in proteins can serve as sensitive markers for changes occurring in the tertiary structure of the proteins during unfolding and aggregation [26, 25, 57, 58, 59, 60].
It is noteworthy that, in the spectral range of 1000-1800 $\mathrm{cm^{-1}}$ (Fig. 4 a), the Raman signal from the water solvent is practically negligible compared to the vibrational signals attributed to the Fib protein. Conversely, the intense spectral distribution characteristic of the OH stretching profile of water entirely dominates the high frequency range between 2800 and 3800 $\mathrm{cm^{-1}}$ in the UVRR spectrum of Fib.
Remarkable spectral variations can be detected as a function of temperature in both explored wavenumber regions for the protein Fib alone (Fig. 4a). In particular, the intensity of the Trp Raman signals W16, W7, and W3 significantly decreases starting from about 50-60C. For the intense signal at $\sim$ 1554 $\mathrm{cm^{-1}}$ (Trp W3), the change is even more noticeable, as the peak almost disappears at high temperatures. Conversely, the signal assigned to the Tyr band at $\sim$ 1616 $\mathrm{cm^{-1}}$ appears stable as a function of temperature, while a new broad spectral contribution centered at about 1650 $\mathrm{cm^{-1}}$ tends to rise with increasing temperature. At the same time, we observe a considerable increase in the total intensity of the OH stretching band of water when the temperature exceeds 50-60C.
The spectra for Fib with CNPs (Fig. 4 b) clearly show that the Raman signal of the Fib protein, when introduced into an NP solution, decreases significantly, and some spectral features almost completely disappear or exhibit broadening. For example, this is evident from the comparison between the spectra of Fib alone and with CNPs in the range of 1500-1800 $\mathrm{cm^{-1}}$. This is partly related to the increase in UV absorption at 220-230 nm of the solution with CNPs compared to the pristine protein. In fact, self-absorption reduces the observed Raman intensities of Fib, which is clearly evident when looking at the total intensity of the OH stretching band of water (Fig. 4 b). More importantly, we can observe that by increasing the temperature, the Raman spectra of Fib with CNPs do not exhibit any remarkable change in both the explored wavenumber ranges, unlike the case of pristine protein. This experimental finding suggests that the presence of CNPs in solution tends to mitigate the effect of temperature on the structure of Fib.
For the spectra of Fib with SiO2NPs, we observe a $T$-dependence of the OH stretching band of water that resembles that of free Fib (Fig. 4 c). A more quantitative insight can be extracted from analyzing the temperature dependence of the intensities of some characteristic Raman signals (Fig. 5).
The integrated area of the OH stretching band for Fib alone (Fig. 5 a) exhibits a clear sigmoidal trend as a function of temperature, indicating a transition that occurs at approximately 57C, likely associated with structural transformations for the Fib protein. The sudden increase in this Raman signal above 50C likely marks the unfolding and aggregation of the protein driven by the temperature rise. The sigmoidal trend suggests that protein denaturation begins at around 50C and rapidly increases until 70C, at which point the intensity of the OH band area stabilizes. Consequently, aggregates may be forming within this temperature range.
When CNPs are added to Fib’s solution, the OH profile significantly decreases in intensity but retains its shape (Fig. 4 b). This may suggest that water is not able to penetrate the protein corona, likely due to protein unfolding on the surface of the NP. Additionally, the corresponding OH-band integrated area shows little to no dependence on $T$ (Fig. 5 a). This finding suggests that the Fib corona is so strongly associated with particles that even with a significant increase in the energy of the solution due to heating, the proteins remain bound to the CNP. Furthermore, the analysis provides no evidence of aggregation.
When SiO2NPs are in solution with Fib, we observe that the OH-band integrated area is intermediate between the two previous cases, with a $T$-dependence that is weaker than for free Fib (Fig. 5 a). Furthermore, the increase in the area occurs at a temperature close to that of the free Fib. This behavior suggests that Fib absorbed onto SiO2NPs undergoes unfolding and aggregation similar to free Fib. Nevertheless, the analysis does not allow us to quantify the amount of unfolding and aggregation relative to the bulk case.
We also analyze the temperature dependence of the integrated area of some characteristic Raman signals in the fingerprint region 1000-1800 $\mathrm{cm^{-1}}$ assigned to Fib (Fig. 5 c-e). In particular, the unfolding and aggregation of proteins should alter the local packing of side chains and the solvent exposure of the residues, affecting the surrounding environment of Trp amino acids and the resonance Raman cross-section of the associated bands [54, 61, 62]. Therefore, by comparing their relative intensity for Fib in the absence and presence of NPs, we can qualitatively assess the effect of NPs on the $T$-dependence of the bands (Fig. 5 c-e).
Interestingly, for bulk Fib, we observe a rise in the spectral contribution between 1660-1670 $\mathrm{cm^{-1}}$ starting from about 55C (Fig. 5 c) and a decrease in the intensity of the Trp W3 band (Fig. 5 d) at about 52C, while unclear features are evident for other peak areas. The two changes could represent two different events, such as unfolding and aggregation, or a single event in which both unfolding and aggregation occur at the same temperature, as debated in the literature. The experimental indeterminacy does not allow us to clearly contribute to the debate, although there are theoretical reasons to support the first interpretation [63].
When the solution includes the NPs, none of the band intensities exhibit a consistent tendency, hindering any clear interpretation (Fig. 5 b-e). Overall, these spectral data reveal the difficulty of interpreting the experimental data and the necessity for a systematic method for their analysis.
2.1.3 Circular Dichroism and UV Absorption Spectroscopy
Circular dichroism (CD) is a widely used technique in biochemistry for studying the secondary structure of biomolecules. By examining the ellipticity $\mathrm{\theta_{\lambda}}$, which is proportional to the differential absorption of circularly polarized light, one can characterize how the chiral structures of proteins, such as alpha helices and beta barrels, change under various conditions. At the same time, the UV absorbance spectra serve as a complementary tool providing important information about protein stability, unfolding, and aggregation.
By combining CD/UV absorption spectroscopy with UVRR experiments, we aim to provide a comprehensive view of the structural changes in the Fib protein’s secondary and tertiary structures while overcoming the intrinsic limitations of each of the three techniques. Therefore, we collect UV and CD spectra as a function of temperature for the Fib protein, both alone and in the presence of NPs (Fig. 6 a-f).
We observe that at approximately 52C in both the CD/UV absorbance spectra of free Fib and Fib with SiO2NPs, there is a noticeable modification of the experimental profiles. Specifically, they indicate i) a decrease in UV absorption (Fig. 6 a and e), and ii) a change in Fib’s secondary structure (Fig. 6 b and f of the CD spectra). In particular, the flattening of the CD spectra indicates a loss of ellipticity, which is sensitive to the protein’s secondary structure, and the decrease in intensity at 208 nm and 222 nm suggests a loss of $\alpha$-helical moieties. These changes correspond to an increase in the intensity of the water band in the Raman spectra for these two cases (Fig. 4 a and c).
On the other hand, the UV absorption of Fib with CNPs shows no $T$-dependence (Fig. 6 c), while the CD spectra, which are noisy due to CNP’s absorption, exhibit a $T$-dependence that is weaker than those observed for free Fib and Fib with SiO2NPs (Fig. 6 d). These data are consistent with the corresponding results from UVRR spectra, displaying a very weak dependence of the spectra on temperature (Fig. 4 b).
These trends can be emphasized by selecting the signals at a specific wavelength and analyzing them as a function of temperature. For UV absorption, we select the signal at 225 nm (Fig. 6 g), while for CD, we choose the ellipticity $\mathrm{\theta\lambda}$ at 222 nm (Fig. 6 h). For both quantities, we observe some similarities between Fib in PBS and Fib with SiO2NPs, while no or weak $T$-dependence for Fib with CNPs.
Therefore, as with the UVRR spectra, while we observe differences and similarities among the three cases, no quantitative results can be systematically extracted from these data, as further discussed in Section 3 (Fig. S3) of the ESI. To achieve an analysis of the spectroscopic data that quantitatively evaluates how far the NP-adsorbed proteins are from their native state, we have developed a method based on an ML approach for the combined spectra.
To prepare the acquired spectral data for treatment using our proposed ML workflow, we first apply data curation through baseline corrections, followed by appropriate spectral truncations, removing the silent regions from the UVRR spectra and the noisy regions from the CD and UV absorption spectra (see Section B.4 for details). The preprocessed spectra for each system (free Fib, Fib with carbon NPs, Fib with silica NPs) are then combined into a higher-dimensional vector that captures all significant chemical information at each temperature.
Next, we implement data standardization, which is typically necessary for PCA and t-SNE analysis. Specifically, this procedure accounts for the fact that the data at hand are measured in different units [34, 38, 64].
Finally, we apply the ML model, which includes (i) metric analysis, (ii) clustering, manifold reduction, and (iii) the subsequent elaboration of the outputs. Different tools, such as Euclidean and Manhattan distances, along with PCA, are routinely used to analyze high-dimensional and spectral data. For example, the distances can measure the similarity between pairs of spectra to quantify the structural changes of proteins. At the same time, PCA can visualize the patterns of the original data spectra in a two-dimensional space. However, these traditional tools encounter issues due to the variety and high dimensionality of the data at hand. For this reason, we replace them with the Wasserstein distance and t-SNE, respectively, as discussed in the next section. Our entire work pipeline, starting from the experiments and concluding with the ML analysis, is illustrated in Fig. 7.
Figure 7: Work pipeline including the automated ML workflow. The first two steps (circles) represent the experiment execution and output spectra. In the case considered here, we prepare solutions containing Fib—either alone or in combination with CNPs or SiO2NPs—and perform spectral measurements of these systems at various temperatures. Next, data curation and standardization of the experimental outputs are applied. Then, preprocessed data are analyzed using metrics computation and manifold learning via unsupervised ML tools, enabling quantitative and qualitative analysis.
3 ML model
The key idea behind our ML model is to use a physics-informed metric to identify and quantify potential protein conformational changes in our solutions, caused by increasing the temperature. However, since the spectral data is high-dimensional, it is also essential to visualize emerging patterns and assess their consistency with the metric analysis. With this in mind, we select the best state-of-the-art tools for this problem: the Wasserstein distance and t-SNE.
The Wasserstein (or Earth Mover’s) distance, defined in Eq. 2, measures the distance between two probability distributions. It circumvents the challenges of the curse of dimensionality by calculating the minimum transport-based cost required to transform one distribution into another, primarily comparing spectra based on their shapes rather than directly relying on spectral intensities [65]. In contrast, similarity measures such as Euclidean and Manhattan distances (defined in Table 2), which are commonly used, focus solely on intensity differences. As a result, they are unreliable when handling high-dimensional data.
The t-SNE is a nonlinear dimensionality reduction technique that converts high-dimensional data into a probability distribution reflecting the similarities among data points. It is specifically designed to address the crowding problem, which refers to the risk of losing local structure and compressing distances when high-dimensional data is mapped to a lower-dimensional space. The t-SNE mitigates the crowding problem because the heavy tails of the Student-t distribution in the low-dimensional space allow for better separation of points that are moderately distant. When applied to complex, high-dimensional data where local relationships are crucial, t-SNE outperforms the faster PCA, which cannot effectively handle data from different sources due to its reliance on variance maximization. This maximization can be significantly influenced by outliers, resulting in misleading results.
In our metric computations, we consider as reference spectra those taken at $T_{0}=22^{\circ}$, which is the lowest temperature of our measurements. At $T_{0}$, Fib is in its native state. As a result, we compute the similarity distance between the composite spectrum at $T_{0}$ and the other spectra at temperatures $T$, separated by $\Delta T=4^{\circ}$. Accordingly, the Wasserstein distance quantifies the increase of the protein’s structural disorder with temperature. To emphasize the change, we normalize the Wasserstein distance with the two values at the two extreme temperatures for the free Fib case (Fig. 8).
We find that the disorder content of the protein increases monotonically with temperature for all systems. In particular, the normalized disorder content for the free Fib and the Fib with silica NPs depends on temperature in a similar way. In both cases, around $(52\pm 2)^{\circ}$, the normalized Wasserstein distance experiences a significant change, while outside that range, the increase is approximately linear.
Comparing this analysis with the original experimental spectra and our knowledge of the temperature range for unfolding the free Fib in solution, we conclude that the characteristic temperature $(52\pm 2)^{\circ}$ indicates the unfolding transition for both free Fib and the Fib with silica NPs. Therefore, the adsorption of Fib onto SiO2NPs does not affect the Fib unfolding temperature.
However, while the free Fib attains the reference value of one at $T=86^{\circ}$, the protein with the silica NP only reaches a value of 0.9 at the same $T$. Hence, the silica surface partially limits the unfolding of Fib.
In contrast, the effect of the carbon interface is significantly different. In this case, the increase in the normalized Wasserstein distance is approximately linear across the entire range of explored temperatures, with no indication of a rapid structural change as in a folding-unfolding transition. Furthermore, the distance at $86^{\circ}$ only reaches a value of 0.5. Conventionally, a value of 0.5 for a folding-unfolding structural parameter is the limit between the two states. Furthermore, for the other two cases, the larger structural change occurs when the Wasserstein parameter is approximately 0.5.
Therefore, the impact of the NP surface is substantial, inducing only a gradual structural change. This analysis enables us to interpret the original spectra better, clarifying that the structural disorder caused by the increase in temperature in this case does not resemble the standard unfolding that occurs for the free protein.
These observations regarding the differences in the Wasserstein distances of Fib proteins interacting with silica NPs and carbon NPs indicate that the two interfaces impact Fib’s structure in fundamentally different ways. Furthermore, recalling that we normalize the Wasserstein distance relative to the minimum and maximum disorder in the free Fib case, Fig. 8 provides information about the absolute value of disorder as a function of $T$ in each case. In particular, Fib on silica NPs unfolds approximately 90% of the free case, while Fib on carbon NPs is about 50% folded at the same $T$.
It is worth noting that using Euclidean and Manhattan distances would result in a different erroneous scenario, where several substantial changes, even at the lowest temperatures, appear across the temperature range. This is discussed in Section 1 of ESI.
Figure 8: Normalized Wasserstein distance computed for composite spectra of free Fib and Fib with NPs. Upper panel: free Fib (squares). Central panel: Fib with CNPs (circles). Lower panel: Fib with SiO2NP (triangles). In each case, the quantity is normalized with respect to the values at $22^{\circ}$ and $86^{\circ}$ for the free case, corresponding to Fib in its native (folded) or disordered (unfolded) state. Therefore, it quantifies the relative distance of the spectra (structural disorder) from the reference spectrum of a free Fib protein in its native state at $22^{\circ}$.
To better interpret the absolute amount of disorder the Fib acquires with temperature increase in each case, we apply t-SNE to the data under scrutiny. This manifold learning algorithm models similarities between composite spectra in the original high-dimensional space using a Gaussian distribution and maps it into a Student-t distribution in the low-dimensional space [32]. It then minimizes the Kullback–Leibler (KL) divergence between the probability distribution of the original high-dimensional data and the low-dimensional space, which in our case is two, visualizing a corresponding two-dimensional embedding [32]. The KL divergence is a non-symmetric measure of similarity between two distributions, which, when minimized, preserves the local structure of the data. Hence, if two points are close in high-dimensional space, a minimum KL divergence keeps them close in the low-dimensional space as well.
To compute t-SNE, an input parameter called the perplexity $\mathcal{P}$ is required. As discussed in Section 4 of ESI, the perplexity is set to 7, consistent with the small size of the datasets. Other parameters are set to their default values for t-SNE optimization [33].
a) Refer to caption
b) Refer to caption
Figure 9: t-SNE analysis of composite spectra. a) Two-dimensional embedding from t-SNE with cosine similarity of composite spectra of free Fib (squares), Fib with CNPs (circles), and Fib with SiO2NPs (triangles), each corresponding to a given temperature between $22^{\circ}$ and $86^{\circ}$, at $4^{\circ}$ intervals. The temperature of each data point is color-coded as indicated in the vertical color bar. . b) Same as in panel a), with centroids (red cross marks) computed using the K-means clustering algorithm, with average silhouette score $\approx 0.94$. In both panels, we use $\mathcal{P}=7$, optimal number of clusters $n_{\rm op}=5$, and PCA initialization as discussed in Section 4 of ESI.
To calculate the embedding, we compute t-SNE using cosine similarity (Eq. 1), as it is expected to make t-SNE less susceptible to the high dimensionality of the spectral data. We identify five distinct clusters (Figure 9a).
All data corresponding to free Fib (squares) and Fib with SiO2NPs (triangles) are divided into two clusters, each characterized by distinct temperature ranges: $T\leq 50^{\circ}$ and $T\geq 54^{\circ}$. The emergence of two clusters for free Fib and Fib with SiO2NPs indicates a significant conformational change around $(52\pm 2)^{\circ}$, consistent with our interpretation of the normalized Wasserstein metrics as a measure of protein disorder, which signifies the folding-unfolding transition of Fib as the temperature increases.
The remaining data for Fib with CNPs (circles) are grouped into a single cluster. This indicates that, at all analyzed temperatures, no significant structural change occurs, further supporting our rationale that a normalized Wasserstein distance with a value of approximately 0.5 corresponds to a protein that is around 50% folded.
The calculation of centroids based on the K-means algorithm, with the number of clusters set to 5, confirms that the data instances are well clustered. The clustering pattern has an average silhouette score of approximately 0.94, indicating strong clustering [64] (Fig. 9 b).
This analysis demonstrates that, for normalized Wasserstein distance up to 0.5, the change in the Fib structure is limited, as with CNPs. Above the 0.5 value, the Wasserstein parameter marks an unfolded state, as with free Fib and Fib with silica NPs.
Although not quantitative, the t-SNE analysis qualitatively confirms that the structural change for Fib with SiO2NPs is less significant than in the free Fib case, even though the Fib unfolds at the same temperature in both scenarios. In contrast, it highlights the gradual increase of disorder without unfolding for Fib adsorbed on the carbon NPs up to $86^{\circ}$.
We emphasize that if PCA is used instead of t-SNE, the resulting two-dimensional projection would suggest the presence of five or more clusters. Furthermore, the calculation of the centroids would be less precise, making it more difficult to quantitatively evaluate the absolute amount of structural disorder in the three cases. This is because PCA identifies some data points as outliers, as discussed in Section 2 of the ESI, due to its difficulty in handling data originating from very different sources.
Finally, we note that a similar embedding is obtained when computing t-SNE with the Euclidean norm/distance and $\mathcal{P}=7$, as discussed in Section 4 of the ESI. In this case, the clustering quality decreases, with the average silhouette score dropping to approximately $0.91$. Nevertheless, five well-defined clusters still emerge, reinforcing the idea that t-SNE is not significantly affected by the data’s high dimensionality.
4 Conlcusion
We developed a method that leverages ML to interpret spectral data from multiple experimental techniques, measuring proteins’ structural properties. This approach optimizes the combined information and systematically evaluates proteins’ relative and absolute structural disorder as a function of varying external conditions.
[…]

### v5

@@ 1 Introduction
The evolution of the protein corona has been described by theory [15] and confirmed experimentally [16, 17], and its dynamic changes are crucial for nanotoxicology and nanomedicine [18]. However, the way the NP surface influences the protein structural changes within the corona remains debated [19, 20, 21, 22]. For instance, theory predicts that proteins near a hydrophobic interface should unfold and aggregate less than those in bulk, and that increasing temperature does not facilitate unfolding and reduces aggregation [23]. Conversely, on a hydrophilic NP, the prediction is that proteins adsorb while remaining folded, and as temperature increases, they unfold but do not desorb [24].
Recent experiments have confirmed that the surface chemistry of NP strongly influences corona formation and protein response. Hydrophobic surfaces promote hemolysis and platelet adhesion [25, 5], while hydrophilic surfaces allow for more tunable coronas [26]. Experiments using NPs with different ligand chemistries [27, 28] and various sizes [29] further show how polarity and curvature affect protein adsorption and secondary structure.
Machine learning (ML) has emerged as a powerful tool for uncovering patterns in complex biological datasets, particularly where subtle spectral differences are difficult to detect visually. For instance, recent research demonstrates that ML models can accurately classify different stages of anesthesia based on near-infrared spectroscopy data [30].
Here, we combine experimental techniques with machine learning (ML) analysis to investigate how Fib undergoes structural changes across various temperatures when exposed to carbon (CNP) and silicon dioxide (SiO2) NPs. We focus on NPs of similar size but with different surface chemistries to isolate the effects of hydrophobicity and hydrophilicity on protein behavior.
To probe protein structural changes, we employ a combination of spectroscopic techniques: UltraViolet Resonance Raman (UVRR) spectroscopy [31], Circular Dichroism (CD) [32], and UV absorption spectroscopy [33]. Although UVRR, CD, and UV absorption provide highly specific spectral fingerprints, differences between conditions can be subtle and not easily visible to the naked eye. Traditional workflows that depend on visual comparison and manual peak fitting or deconvolution are operator-dependent and can yield inconsistent results, especially when spectra are noisy or have overlapping features. Furthermore, modern instruments regularly generate large, multi-modal datasets, making thorough manual analysis impractical. These limitations motivate our unsupervised ML approach, which enables us to extract quantitative, reproducible descriptors of conformational change from complex, multi-modal datasets, while minimizing operator bias and the constraints of manual spectral analysis. In fact, we chose unsupervised ML because there are no reliable labels for conformational states of our experimental system in these spectra, and supervised models would risk bias.
We examine high-dimensional, multi-component spectroscopic data using a novel unsupervised ML workflow. This approach is designed to quantify and visualize protein conformational changes as temperature varies. Spectroscopic datasets often contain subtle structural transitions embedded within complex, high-dimensional data that are affected by noise and collinearity. These data typically include hundreds to thousands of correlated variables, characterized by information redundancy, small sample sizes, and significant uncertainty, resulting in the curse of dimensionality, which hinders effective analysis and interpretation [34].
Noise arising from instrumental limitations, environmental fluctuations, and sample variability can distort the data’s intrinsic structure, reducing the effectiveness of traditional dimensionality reduction techniques [35]. Additionally, collinearity among spectral bands inflates the feature space without adding meaningful information, increasing the risk of overfitting and limiting model generalizability [36]. To address the computational challenges of integrating multispectral data, as described above, advanced ML methods are essential for uncovering the underlying low-dimensional structure of the data. Manifold learning techniques like t-distributed stochastic neighbor embedding (t-SNE) have proven to be powerful tools, providing nonlinear dimensionality reduction capabilities that preserve both local and global data relationships, along with computational efficiency [37]. The novelty of our proposed ML model rests on two key algorithms: the noisy-resistant Wasserstein distance [38, 39] and t-SNE [40, 41, 42].
Our analysis reveals distinct effects of NP surface chemistry on protein structural disorder and aggregation. Fib adsorbed onto CNPs maintains a partially unfolded structure that is only weakly dependent on temperature, while Fib on SiO2NPs exhibits a temperature-dependent increase in disorder similar to free Fib, with a marked transition around 50C. These findings, supported by unsupervised ML analysis, provide experimental confirmation of theoretical predictions and underscore the importance of NP surface properties in protein adsorption and stability.
2 Materials and Methods
2.1 Materials
We use monodisperse nanoparticles of similar size but differing surface chemistry. SiO2NPs of 0.1 µm (stock concentration of 50 mg/mL) were provided by Kisker Biotech GmbH (Germany). CNPs were synthesized in the lab. CNPs and SiO2NP have a comparable geometrical and hydrodynamic diameter to prevent any effects caused by a differing surface curvature.
Both exhibit a negative zeta potential at neutral pH [5, 43], indicating the presence of weakly acidic groups on the surface. In the case of CNPs, these species include acidic carboxylic or phenolic groups formed during synthesis, while for SiO2NP, they consist of surface hydroxyl groups (silanols, Si-OH).
The density of acidic groups on the NP surfaces is expected to be in the same range for both samples (around 4-6 groups per nm2) [5]. However, the two surfaces exhibit significant differences. While in CNPs the oxygenated moieties are bound to a hydrophobic surface composed of C-C bonds, in the case of silica, silanols are attached to a more hydrophilic framework made up of polar Si-O-Si bridges. Therefore, while SiO2NPs are hydrophilic, CNPs are more hydrophobic.
[…]